OpenStudy (anonymous):

3(the square root of) a minus 5 = 13. The teacher said the correct answer was a = 36. I've done similar problems before but I can't figure out how to get a = 36... I did the normal steps, square both sides, etc. Can anyone help me?

OpenStudy (anonymous):

\[3\sqrt{a}-5\]

OpenStudy (anonymous):

\[\sqrt[3]{x-5}=13\] or \[3\sqrt{x-5}=13\]

OpenStudy (anonymous):

or \[3\sqrt{x}-5=13\]

OpenStudy (anonymous):

The second one

OpenStudy (anonymous):

3rd example you wrote, second comment

OpenStudy (anonymous):

\[3\sqrt{x}-5=13\] add 5 and get \[3\sqrt{x}=18\] divide by \(3\) to get \[\sqrt{x}=6\]

OpenStudy (anonymous):

then square

OpenStudy (anonymous):

Thanks so much! I was trying to square first and it was making the problem really complicated.

OpenStudy (anonymous):

yeah that will be a pain isolate the radical first if you can

OpenStudy (anonymous):

Will do thanks for advice an help!